### v1

# supplement.pdf

Abstract
Random fluctuations caused by environmental noise can lead to decoherence in quantum systems. Exploring and controlling such dissipative processes is both fundamentally intriguing and essential for harnessing quantum systems to achieve practical advantages and deeper insights. In this Letter, we first demonstrate the diffusive dynamics assisted by controlled dephasing noise in superconducting quantum circuits, contrasting with coherent evolution. We show that dephasing can enhance localization in a superconducting qubit array with quasiperiodic order, even in the regime where all eigenstates remain spatially extended for the coherent counterpart. Furthermore, by preparing different excitation distributions in the qubit array, we observe that a more localized initial state relaxes to a uniformly distributed mixed state faster with dephasing noise, illustrating another counterintuitive phenomenon called Mpemba effect, i.e., a far-from-equilibrium state can relax toward the equilibrium faster. These results deepen our understanding of diffusive dynamics at the microscopic level, and demonstrate controlled dissipative processes as a valuable tool for investigating Markovian open quantum systems.
Decoherence is an inherent feature of quantum systems, arising from the inevitable interactions with the numerous degrees of freedom in the surrounding environment. Such random interactions generally lead to two types of dissipative processes in quantum systems, i.e., energy relaxation caused by spontaneous emission, and dephasing caused by random fluctuations in the system parameters. Exploring and controlling these dissipative processes is not only interesting in itself from fundamental aspects, but also crucial for harnessing quantum systems to gain any technical advantage or insight Harrington et al. (2022). Uncontrolled dissipative processes pose a major challenge in quantum information processing, as they erode essential quantum resources like coherence and entanglement, degrade the fidelity of quantum gates, and add noise to measurement signals. However, engineered dissipative processes can play a constructive role, for example, in quantum measurement Clerk et al. (2010); Chen et al. (2024a), quantum state stabilization Lin et al. (2013); Leghtas et al. (2015); Lu et al. (2017); Grimm et al. (2020); Lingenfelter et al. (2024), autonomous quantum error correction Gertler et al. (2021); Xu et al. (2023); Li et al. (2024a), dissipative phase transition Prosen and Pižorn (2008); Kessler et al. (2012); Roberts and Clerk (2023a, b), long-range coherence Dutta and Cooper (2020); Li et al. (2023), and non-Hermitian systems Bergholtz et al. (2021).
While dephasing noise generally spoils quantum correlations, it can lead to collective phenomena, including noise-induced transport Lindner and Schimansky-Geier (2002); Gopalakrishnan et al. (2017); Ren et al. (2020), quantum synchronization Schmolke and Lutz (2022); Tao et al. (2024), and quantum phase transition Liu et al. (2018). In quantum many-body systems, dephasing noise acts as a dynamically fluctuating disorder, which is generally acknowledged to enhance transport and spoil localization Gurvitz (2000); Plenio and Huelga (2008); Levi et al. (2012); Nakajima et al. (2018). Recently, this common wisdom has been theoretically revisited, finding that localization can be enhanced by dephasing noise in one-dimensional (1D) quasiperiodic systems, even in the regime where all the eigenstates are spatially extended in the coherent counterpart Longhi (2024a), as illustrated in Fig. 1(a). The excitation in the lattice spreads via a diffusive process assisted by the dephasing effect, showing distinct localization behavior in contrast with the ballistic dynamics under coherent evolution. Dephasing noise can also assist the illustration of another counterintuitive phenomenon called Mpemba effect in the diffusive dynamics Longhi (2024b), which means a far-from-equilibrium state can relax toward equilibrium faster Bechhoefer et al. (2021), as illustrated in Fig. 1(b). These theoretical findings provided rigorous insights into the quantum dynamics besides coherent evolution Carollo et al. (2021), appealing for further experimental explorations.
Figure 1: Dephasing-assisted transition from ballistic to diffusive transport and accelerated equilibration in a 1D qubit array. (a) Illustration of the density profile for an initial state $|\psi_{0}\rangle$ evolved under Hamiltonian $H$ without (upper panel) and with dephasing noise (lower panel). (b) Schematic for Mpemba effect, where a far-from-equilibrium state reaches equilibrium faster than a state closer to equilibrium under certain conditions (here with dephasing noise). (c) The 1D array model used in this experiment to demonstrate the above phenomena, where controlled dephasing is incorporated by randomly modulating the qubit frequencies on each site.
Featuring scalability and controllability, superconducting quantum circuits are not only one of the leading candidates for quantum computation Acharya et al. (2023), but also an ideal playground for exploring exotic phenomena with controlled dissipation. For example, superconducting qubits have been used to explore dissipative many-body dynamics Ma et al. (2019); Mi et al. (2024); Du et al. (2024), non-Hermitian physics Naghiloo et al. (2019); Chen et al. (2021, 2022), and quantum state stabilization Touzard et al. (2018); Li et al. (2024b); Brown et al. (2022); Chen et al. (2024b); Li et al. (2024c). In this Letter, we demonstrate the dephasing-assisted diffusive dynamics and associated phenomena using a 1D array of seven superconducting qubits with tunable nearest-neighbor couplings, as illustrated in Fig. 1(c). The dynamics of this system are captured by the effective Hamiltonian of a 1D tight-binding model with $L=2l+1$ sites:
$H/\hbar=\sum_{j=-l}^{l}\Delta\omega_{j}\sigma_{j}^{+}\sigma_{j}^{-}+\sum_{j=-l+1}^{l}g_{j}(\sigma_{j}^{+}\sigma_{j-1}^{-}+\sigma_{j}^{-}\sigma_{j-1}^{+}),$ (1)
where $\sigma^{+}_{j}$ ($\sigma_{j}^{-}$) represents the raising (lowering) operator for the qubit at site $j$, $\sigma_{j}^{+}|0\rangle^{\otimes L}=|1_{j}\rangle$, $\Delta\omega_{j}/2\pi$ is the frequency detuning of the qubit relative to the average system frequency, $g_{j}/2\pi$ is the tunable coupling strength between site $j-1$ and $j$, and $l=3$ here, see Supplemental Material for more details of the device. By adjusting the qubit Z bias, we can tune the qubits on resonance with $\Delta\omega_{j}=0$. To introduce controlled dephasing noise into the system, we apply a series of white noises $\xi_{j}(t)$ to the modulation of the qubit frequency $\Delta\omega_{j}$. The evolution of density matrix $\rho$ can be effectively described by the Markovian dynamics generated in the Lindblad master equation:
$\frac{d\rho}{dt}=-\frac{i}{\hbar}\left[H,\rho\right]+\sum_{j}\left(\mathcal{L}_{j}\rho\mathcal{L}_{j}^{\dagger}-\frac{1}{2}\left\{\mathcal{L}_{j}^{\dagger}\mathcal{L}_{j},\rho\right\}\right),$ (2)
where $\mathcal{L}_{j}=\sqrt{\Gamma_{j}}\sigma_{j}^{+}\sigma_{j}^{-}$ is the collapse operator for dephasing, $\Gamma_{j}$ is the effective dephasing rate introduced by the noise $\xi_{j}(t)$ with $\Gamma_{j}=\Gamma$, see Supplemental Material for details.
Figure 2: Diffusive spreading dynamics with dephasing noise. (a) Schematic of the experimental pulse sequence. (b)-(c) Measured population dynamics $n_{j}$ without (b) and with (c) dephasing noise, illustrating ballistic- and diffusive-like spreading, respectively. (d) The integrated moment $M(t)$ as a function of normalized time $Jt$, which captures the ballistic (b) and diffusive manner of spreading (c) during early dynamics. Dots indicate measured results, while the red (blue) solid line denotes the scaling function of $M\sim t^{1}$ ($t^{0.5}$).
Figure 3: Dephasing-enhanced localization in a quasiperiodic lattice. (a) Schematic for the off-diagonal Aubry-André model with aperiodic coupling strengths, where the controlled dephasing noises are injected into the qubits. (b) Numerical calculation of $M$ versus $\kappa$ with and without dephasing noise for $L=89$ sites. (c)-(f) Measured population dynamics $n_{j}(t)$ without (c,e) and with (d,f) noise, showing oscillatory dynamics and diffusive spreading behavior, respectively. (g)-(h) The integrated moment $M$ evaluated from the measured dynamics shown in (c-f), where the solid bars and black frames denote the measured and numerical results. The ratio $\kappa=B/A$ is $0.3$ for (c,d,g), and $0.7$ for (e,f,h).
[…]
Dephasing-enhanced localization.– Recently, the off-diagonal Aubry-André model Aubry and André (1980); Cestari et al. (2016); Xiao et al. (2021) has been revisited with dephasing noise Longhi (2024a), finding that the spreading dynamics of excitation can exhibit a different localization behavior from the coherent evolution. We consider this model shown in Fig. 3(a), where the incommensurate aperiodic coupling strength is introduced as $g_{j}=A+B\cos(2\pi\alpha j+\theta)$, with $A,B>0$, irrational frequency $\alpha=\left(\sqrt{5}-1\right)/2$, and phase offset $\theta=0$ throughout this work. In the regime of $\kappa=B/A<1$ for the off-diagonal Aubry-André model, all the eigenstates are spatially extended Xiao et al. (2021). In Fig. 3(b), we numerically show the behavior of the integrated moment $M$ for different values of $\kappa$, both with and without dephasing noise, where $L=89$ is a Fibonacci number. It is evident that $M$ without dephasing noise is an order of magnitude larger than that with noise. To experimentally investigate the phenomena of localization, we focus on the early time dynamics of a finite system tailored to the superconducting quantum circuits, in which localization behavior can still be distinguished from the experimental observables Xiao et al. (2021); Bravyi et al. (2006).
Here, we present the experimental demonstration for dephasing-enhanced localization described by the off-diagonal Aubry-André model with dephasing noise. By configuring the coupling strength as quasiperiodic order through the tunable couplers between qubits, we investigate such intriguing phenomena by the characteristic dynamics of qubit excitation. We first explore the coherent evolution without artificial dephasing noise, where $A+B=J\approx 2\pi\times 8.3~{}\mathrm{MHz}$, and different $\kappa$ within the regime of extended phase. We initialize a single-particle excitation in the center site, and observe the population of each site versus normalized time $Jt$ as shown in Fig. 3(c) and (e) for $\kappa=0.3$ and 0.7, respectively, where the dynamics show clear oscillations traversing all sites, consistent with the picture of extended phase. When the applied noises introduce an intense dephasing noise with $\Gamma/J\approx 30$ on the lattice system, localized eigenstates with an extremely long lifetime emerge, which has a significant impact on the dynamic spreading of excitation, predicting a largely slowing down for the spreading dynamics and an enhanced localization of excitation in the lattice assisted by the dephasing effect Longhi (2024a). The measured dynamics with dephasing noise for $\kappa=0.3$ and 0.7 are shown in Fig. 3(d) and (f) respectively, showing a significantly slowing down diffusive spreading dynamics as compared to their counterpart without noise. The slowing down of diffusive spreading is also captured by the integrated moment $M$ accumulated during the evolution in Fig. 3(d) and (f), as shown in Fig. 3(g) and (h). By comparing the results in Fig. 3(g) and (h), it is clear to see that the dephasing-assisted localization is enhanced when $\kappa$ increases.
Figure 4: Mpemba effect with dephasing noise. (a)-(b) Measured population dynamics $n_{j}$ with a dephasing rate $\Gamma/J\approx 3$, where the system is initially prepared in states with density matrices (a) $\rho_{1}=|1_{0}\rangle\langle 1_{0}|$, and (b) $\rho_{4}=1/4\sum_{j=0}^{3}|1_{j}\rangle\langle 1_{j}|$. (c) Measured distance function $D(t)$ from the equilibrium for initial states $\rho_{1}=|1_{0}\rangle\langle 1_{0}|$, $\rho_{2}=1/2\sum_{j=2}^{3}|1_{j}\rangle\langle 1_{j}|$, $\rho_{3}=1/3\sum_{j=1}^{3}|1_{j}\rangle\langle 1_{j}|$, and $\rho_{4}=1/4\sum_{j=0}^{3}|1_{j}\rangle\langle 1_{j}|$. Solid lines are numerical simulation results. $\rho_{1}$, the farthest from the equilibrium state, relaxes faster than the other three mixed states closer to the equilibrium state $\rho_{E}$, illustrating the Mpemba effect.
Mpemba effect with dephasing noise.– We finally explore the Mpemba effect from the dynamical behavior toward the equilibrium of the diffusive dynamics. The so-called Mpemba effect is unveiled by quantitatively evaluating the relaxation dynamics Longhi (2024b), which describes a scenario where a far-from-equilibrium state can relax toward equilibrium faster than a state closer to equilibrium Bechhoefer et al. (2021). The classic example is that warmer water can sometimes freeze faster than colder water Mpemba and Osborne (1969). Recently, quantum versions of this effect Carollo et al. (2021) have been investigated in different quantum systems including a quantum dot coupled with two reservoirs Chatterjee et al. (2023), coherently driven trapped ion qubit coupled to a thermal Markovian bath Aharony Shapira et al. (2024), and an array of trapped ions with power-law decaying interactions Joshi et al. (2024). In the single-excitation subspace of the 1D system where the dephasing rate is much larger than energy relaxation, any initial state has a non-vanishing projection onto the equilibrium state $\rho_{E}=1/L\sum_{j}|1_{j}\rangle\langle 1_{j}|$ spreading in all sites of the lattice, which leads to a relaxation dynamics toward $\rho_{E}$ in a long-time limit $Rt\gg 1$ with a diffusion rate $R=2J^{2}/\Gamma$ Longhi (2024a). The Mpemba effect with dephasing noise is evaluated by the observable distance function $D(t)$ for a given state $\rho(t)$ from the equilibrium state $\rho_{E}$ defined as $D(t)=\operatorname{Tr}\left[\rho(t)\log\rho(t)\right]-\operatorname{Tr}\left[\rho_{E}(t)\log\rho_{E}(t)\right]$. In the case of single-particle excitation, it takes the following form in terms of the population $n_{j}(t)$ Longhi (2024b)
$D(t)=\log L+\sum_{j}n_{j}(t)\log n_{j}(t).$ (3)
To experimentally observe the Mpemba effect, a small system size is more favorable Longhi (2024b). As the system size $L$ increases, the uniformly distributed excitation of the equilibrium state has a smaller population at each site, and the relaxation dynamics become slower to reach equilibrium.
We now demonstrate the Mpemba effect in our 1D chain of seven qubits, with the homogeneous coupling strengths $g_{j}=J\approx 2\pi\times 8.3$ MHz and $\Gamma/J\approx 3$ to accelerate the diffusive dynamics toward equilibrium state $\rho_{E}$. As shown in Fig. 4(a), when the system is initially prepared as a highly localized pure state $\rho_{1}=|1_{j=0}\rangle\langle 1_{j=0}|$ with the excitation at site $j=0$, it shows spreading of qubit population $n_{j}(t)$ in all sites when the evolution time $t$ is large enough. In a similar scenario, we initialize the system as a mixed state $\rho_{4}=1/4\sum_{j=0}^{3}|1_{j}\rangle\langle 1_{j}|$ by randomly applying the $\pi$ pulse on the sites $j=0,1,2,3$, which is closer to the stationary state $\rho_{E}$ and gives a temporal evolution in Fig. 4(b), showing a slower diffusion toward the missing occupation on the other sites. Intuitively, the initial state with a closer distance to the equilibrium state is expected to diffuse faster and fill the missing occupation of other sites. To perform a systematic study, we prepare two more initial states $\rho_{2}=1/2\sum_{j=2}^{3}|1_{j}\rangle\langle 1_{j}|$, $\rho_{3}=1/3\sum_{j=1}^{3}|1_{j}\rangle\langle 1_{j}|$ between $\rho_{1}$ and $\rho_{4}$ in terms of distance to $\rho_{E}$, and observe their dynamics towards equilibrium similar to Fig. 4(a) and (b), see Supplemental Material for details and numerical simulations. The deviation between the measured population and corresponding numerical results is mainly limited by two-level system defects and readout errors. By evaluating the distance function $D(t)$ from the measured population $n_{j}(t)$, we present the results of $D(t)$ for different initial states in Fig. 4(c), which experimentally demonstrate that the farthest-from equilibrium state $\rho_{1}$ relaxes faster than the other three mixed states being closer to the equilibrium state $\rho_{E}$, clearly illustrating the Mpemba effect.
Conclusions.– In this work, we experimentally demonstrate the diffusive dynamics assisted by controlled dephasing noise in a 1D array of superconducting qubits. By properly arranging the coupling strengths between neighboring qubits, we implement a quasiperiodic lattice and demonstrate that dephasing can enhance excitation localization in this system, even in the regime where all the eigenstates are spatially extended for the coherent counterpart Longhi (2024a). Moreover, we explore the Mpemba effect in this 1D model by preparing different excitation distributions in the qubit array and observing that more localized initial states relax to a uniformly distributed mixed state more rapidly under dephasing noise Longhi (2024b); Moroder et al. (2024). The experimental demonstration of these two counterexamples not only deepens our understanding of diffusive dynamics at microscopic level, but also highlights the expanded potential of controlled dissipative processes in quantum simulation. Utilizing the dissipative effect synthesized in such fully controlled quantum systems, further efforts can be made for the studies, including exotic non-equilibrium phases Cai et al. (2014); Maghrebi and Gorshkov (2016), reservoir computing approach Angelatos et al. (2021), and quantum thermodynamics Talkner and Hänggi (2020); Cech et al. (2023).
Acknowledgments– We thank Yucheng Wang, Xinchi Zhou, and Yuanzhen Chen for their helpful discussions. This work was supported by the Science, Technology and Innovation Commission of Shenzhen Municipality (KQTD20210811090049034, RCBS20231211090824040, RCBS20231211090815032), the National Natural Science Foundation of China (12174178, 12204228, 12374474 and 123b2071), the Innovation Program for Quantum Science and Technology (2021ZD0301703), the Shenzhen-Hong Kong Cooperation Zone for Technology and Innovation (HZQB-KCZYB-2020050), and Guangdong Basic and Applied Basic Research Foundation (2024A1515011714, 2022A1515110615).