### v2

# Dephasing-assisted diffusive dynamics in superconducting quantum circuits

Abstract
Random fluctuations caused by environmental noise can lead to decoherence in quantum systems. Exploring and controlling such dissipative processes is both fundamentally intriguing and essential for harnessing quantum systems to gain practical advantages and deeper insights. In this work, we first demonstrate the diffusive dynamics assisted by controlled dephasing noise in superconducting quantum circuits, contrasting with coherent evolution. We show that dephasing can give distinct dynamical behavior in a superconducting qubit array with quasiperiodic order. Furthermore, by preparing different excitation distributions in the qubit array, we observe that a more localized initial state relaxes to a uniformly distributed mixed state faster with dephasing noise, illustrating another counterintuitive phenomenon called Mpemba-effect-like quantum dynamics, i.e., a far-from-equilibrium state can relax toward the equilibrium faster. These results deepen our understanding of diffusive dynamics at the microscopic level, and demonstrate controlled dissipative processes as a valuable tool for investigating Markovian open quantum systems.
Decoherence is an inherent feature of quantum systems, arising from the inevitable interactions with the numerous degrees of freedom in the surrounding environment. Such random interactions generally lead to two types of dissipative processes in quantum systems, i.e., energy relaxation caused by spontaneous emission, and dephasing caused by random fluctuations in the system parameters. Exploring and controlling these dissipative processes is not only interesting in itself from fundamental aspects, but also crucial for harnessing quantum systems to gain any technical advantage or insight 1. Uncontrolled dissipative processes pose a major challenge in quantum information processing, as they erode essential quantum resources like coherence and entanglement, degrade the fidelity of quantum gates, and add noise to measurement signals. However, engineered dissipative processes can play a constructive role, for example, in quantum measurement 2; 3, quantum state stabilization 4; 5; 6; 7; 8, autonomous quantum error correction 9; 10; 11, dissipative phase transition 12; 13; 14; 15, long-range coherence 16; 17, and non-Hermitian systems 18.
While dephasing noise generally spoils quantum correlations, it can lead to collective phenomena, including noise-induced transport 19; 20; 21, quantum synchronization 22; 23, and quantum phase transition 24. In quantum many-body systems, dephasing noise acts as a dynamically fluctuating disorder, which is generally acknowledged to enhance transport and spoil localized dynamical behavior 25; 26; 27; 28. Recently, this common wisdom has been theoretically revisited, finding that the dynamical behavior can be changed by dephasing noise in one-dimensional (1D) quasiperiodic systems, in the regime where all the eigenstates are spatially extended in the coherent counterpart 29, as illustrated in Fig. 1(a). The excitation in the lattice spreads via a diffusive process assisted by the dephasing effect, showing distinct dynamical behavior in contrast with the ballistic dynamics under coherent evolution. Dephasing noise can also assist the illustration of another counterintuitive phenomenon called Mpemba-effect-like quantum dynamics 30, which means a far-from-equilibrium state can relax toward equilibrium faster 31, as illustrated in Fig. 1(b). These theoretical findings provided rigorous insights into the quantum dynamics besides coherent evolution 32, appealing for further experimental explorations.
Fig. 1: Dephasing-assisted transition from ballistic to diffusive transport and accelerated equilibration in a 1D qubit array. (a) Illustration of the density profile for an initial state $|\psi_{0}\rangle$ evolved under Hamiltonian $H$ without (upper panel) and with dephasing noise (lower panel). (b) Schematic for Mpemba-effect-like dynamics, where a far-from-equilibrium state (red line) reaches equilibrium (denoted as blue dashed line) faster than a state closer to equilibrium (yellow line) under certain conditions (here with dephasing noise). (c) The 1D array model used in this experiment to demonstrate the above phenomena, where controlled dephasing is incorporated by randomly modulating the qubit frequencies on each site.
Featuring scalability and controllability, superconducting quantum circuits are not only one of the leading candidates for quantum computation 33, but also an ideal playground for exploring exotic phenomena with controlled dissipation. For example, superconducting qubits have been used to explore dissipative many-body dynamics 34; 35; 36, non-Hermitian physics 37; 38; 39, and quantum state stabilization 40; 41; 42; 43; 44. In this work, we demonstrate the dephasing-assisted diffusive dynamics and associated phenomena using a 1D array of seven superconducting qubits with tunable nearest-neighbor couplings, as illustrated in Fig. 1(c). The dynamics of this system are captured by the effective Hamiltonian of a 1D tight-binding model with $L=2l+1$ sites:
$H/\hbar=\sum_{j=-l}^{l}\Delta\omega_{j}\sigma_{j}^{+}\sigma_{j}^{-}+\sum_{j=-l+1}^{l}g_{j}(\sigma_{j}^{+}\sigma_{j-1}^{-}+\sigma_{j}^{-}\sigma_{j-1}^{+}),$ (1)
where $\sigma^{+}_{j}$ ($\sigma_{j}^{-}$) represents the raising (lowering) operator for the qubit at site $j$, $\sigma_{j}^{+}|0\rangle^{\otimes L}=|1_{j}\rangle$, $\Delta\omega_{j}/2\pi$ is the frequency detuning of the qubit relative to the average system frequency, $g_{j}/2\pi$ is the tunable coupling strength between site j1j-1 and $j$, and $l=3$ here, see Supplemental Material for more details of the device. By adjusting the amplitude of microwave pulses applied to the qubits, we can tune the qubits on resonance with $\Delta\omega_{j}=0$. To introduce controlled dephasing noise into the system, we apply a series of white noises $\xi_{j}(t)$ to the modulation of the qubit frequency $\Delta\omega_{j}$. The evolution of density matrix $\rho$ can be effectively described by the Markovian dynamics generated in the Lindblad master equation:
$\frac{d\rho}{dt}=-\frac{i}{\hbar}\left[H,\rho\right]+\sum_{j}\left(\mathcal{L}_{j}\rho\mathcal{L}_{j}^{\dagger}-\frac{1}{2}\left\{\mathcal{L}_{j}^{\dagger}\mathcal{L}_{j},\rho\right\}\right),$ (2)
where $\mathcal{L}_{j}=\sqrt{\Gamma_{j}}\sigma_{j}^{+}\sigma_{j}^{-}$ is the collapse operator for dephasing, $\Gamma_{j}$ is the effective dephasing rate introduced by the noise $\xi_{j}(t)$ with $\Gamma_{j}=\Gamma$, see Supplemental Material for details.
Fig. 2: Diffusive spreading dynamics with dephasing noise. (a) Schematic of the experimental pulse sequence. (b)-(c) Measured population dynamics $n_{j}$ without (b) and with (c) dephasing noise, illustrating ballistic- and diffusive-like spreading, respectively. (d) The integrated moment $M(t)$ as a function of normalized time JtJt, which captures the ballistic (b) and diffusive manner of spreading (c) during early dynamics. Dots indicate measured results, while the red (blue) solid line denotes the scaling function of $M\sim t^{1}$ ($t^{0.5}$).
Fig. 3: Dephasing-assisted dynamics in a quasiperiodic lattice. (a) Schematic for the off-diagonal Aubry-André model with aperiodic coupling strengths, where the controlled dephasing noises are injected into the qubits. (b) Numerical calculation of $M$ versus $\kappa$ with and without dephasing noise for $L=89$ sites. (c)-(f) Measured population dynamics $n_{j}(t)$ without (c,e) and with (d,f) noise for $L=7$ qubits, showing oscillatory dynamics and diffusive spreading behavior, respectively. (g)-(h) The integrated moment $M$ evaluated from the measured dynamics shown in (c-f), where the solid bars and black frames denote the measured and numerical results. The ratio $\kappa=B/A$ is $0.3$ for (c,d,g), and $0.7$ for (e,f,h).
We first demonstrate the transition from ballistic- to diffusive-like dynamics by investigating the qubit evolution. In Fig. 2, we present the temporal evolution of the measured on-site population $n_{j}(t)=\langle 1_{j}|\rho(t)|1_{j}\rangle$ of individual qubits versus normalized evolution time $Jt$. The coupling strengths $g_{j}=J\approx 2\pi\times 8.3~\mathrm{MHz}$ are homogeneous, and the effective reduced dephasing rate $\Gamma/J\approx 30\gg 1$. We examine the population dynamics of the qubit array by initially preparing the central qubit in its excited state $|1_{j=0}\rangle$ using a $\pi$ pulse (see Fig. 2(a)). In the absence of noise, we should observe the excitation spreading in a ballistic manner throughout the lattice. This ballistic-like spreading is confirmed in Fig. 2(b), where the population extends rapidly across the qubits, with reflections at open boundaries due to the finite system size. When noise is applied, the dephasing effect leads to a slowdown in spreading dynamics, transforming its behavior from ballistic to diffusive, as shown in Fig. 2(c). To further analyze the spreading dynamics quantitatively, we use the integrated moment $M(t)=(1/t)\int_{0}^{t}[W(\tau)-W(0)]^{1/2}d\tau$ with the effective second moment $W(\tau)=\sum_{j}{|j-j_{0}|^{2}}n_{j}(\tau)$ and initial center $j_{0}=0$ as shown in Fig. 2(d), which captures the expansion of population dynamics from the initial center, giving the characteristic feature of $M(t)\sim t^{1}$ for ballistic spreading and $M(t)\sim t^{0.5}$ for diffusive spreading with strong dephasing noise. The saturated values of $M$ after $Jt\approx 4$ in the absence of noise are captured from the reflections of qubit population at open boundaries due to the finite system size shown in Fig. 2(b).
Recently, the off-diagonal Aubry-André model 45; 46; 47 has been revisited with dephasing noise 29, finding that the spreading dynamics of excitation can exhibit a different dynamical behavior from the coherent evolution. We consider this model shown in Fig. 3(a), where the incommensurate aperiodic coupling strength is introduced as $g_{j}=A+B\cos(2\pi\alpha j+\theta)$, with $A,B>0$, irrational frequency $\alpha=\left(\sqrt{5}-1\right)/2$, and phase offset $\theta=0$ throughout this work. In the regime of $\kappa=B/A<1$ for the off-diagonal Aubry-André model, all the eigenstates are spatially extended 47. In Fig. 3(b), we numerically show the behavior of the integrated moment $M$ for different values of $\kappa$, both with and without dephasing noise, where $L=89$ is a Fibonacci number used to numerically evaluate the model with a large system size, and the time scale from $Jt=0$ to $Jt=50$ is chosen sufficiently large for numerically evaluating the integrated moment $M$ with the system size $L=89$. It is evident that $M$ without dephasing noise is an order of magnitude larger than that with noise. To experimentally investigate the quantum dynamics, we focus on the early time dynamics of a small finite system tailored to the superconducting quantum circuits, in which the dynamical behavior can still be distinguished from the experimental observables 47; 48.
Here, we present the experimental demonstration for dephasing-assisted dynamics described by the off-diagonal Aubry-André model with dephasing noise. By configuring the coupling strength as quasiperiodic order through the tunable couplers between qubits, we investigate such intriguing phenomena by the characteristic dynamics of qubit excitation. We first explore the coherent evolution without artificial dephasing noise, where $A+B=J\approx 2\pi\times 8.3~\mathrm{MHz}$, and different $\kappa$ within the regime of extended phase. We initialize a single-particle excitation in the center site, and observe the population of each site versus normalized time $Jt$ as shown in Figs. 3(c) and 3(e) for $\kappa=0.3$ and 0.7, respectively, where the dynamics show clear oscillations traversing all sites, consistent with the picture of extended phase. When the applied noises introduce an intense dephasing noise with $\Gamma/J\approx 30$ on the lattice system, localized eigenstates with an extremely long lifetime emerge, which strongly influences the dynamic spreading of excitation and slows down the spreading dynamics 29. The measured dynamics with dephasing noise for $\kappa=0.3$ and 0.7 are shown in Figs. 3(d) and 3(f) respectively, showing a significantly slowing down diffusive spreading dynamics as compared to their counterpart without noise. The slowing down of diffusive spreading is also captured by the integrated moment $M$ accumulated during the evolution from $t=0~\mathrm{ns}$ to $t=400~\mathrm{ns}$ in Figs. 3(d) and 3(f), as shown in Figs. 3(g) and 3(h).
Fig. 4: Mpemba-effect-like dynamics with dephasing noise. (a)-(b) Measured population dynamics $n_{j}$ with a dephasing rate $\Gamma/J\approx 3$, where the system is initially prepared in states with density matrices (a) $\rho_{1}=|1_{0}\rangle\langle 1_{0}|$, and (b) $\rho_{4}=1/4\sum_{j=0}^{3}|1_{j}\rangle\langle 1_{j}|$. (c) Measured distance function $D(t)$ from the equilibrium for initial states $\rho_{1}=|1_{0}\rangle\langle 1_{0}|$, $\rho_{2}=1/2\sum_{j=2}^{3}|1_{j}\rangle\langle 1_{j}|$, $\rho_{3}=1/3\sum_{j=1}^{3}|1_{j}\rangle\langle 1_{j}|$, and $\rho_{4}=1/4\sum_{j=0}^{3}|1_{j}\rangle\langle 1_{j}|$. Solid lines are numerical simulation results. $\rho_{1}$, the farthest from the equilibrium state, relaxes faster than the other three mixed states closer to the equilibrium state $\rho_{E}=1/7\sum_{j}|1_{j}\rangle\langle 1_{j}|$, illustrating the Mpemba-effect-like dynamics.
We finally explore the Mpemba-effect-like dynamics from the dynamical behavior toward the equilibrium of the diffusive dynamics. The so-called Mpemba-effect-like dynamics is unveiled by quantitatively evaluating the relaxation dynamics 30, which describes a scenario where a far-from-equilibrium state can relax toward equilibrium faster than a state closer to equilibrium 31. The classic example is that warmer water can sometimes freeze faster than colder water 49. Recently, quantum versions of this effect 32 have been investigated in different quantum systems, including a quantum dot coupled with two reservoirs 50, a coherently driven trapped ion qubit coupled to a thermal Markovian bath 51, and an array of trapped ions with power-law decaying interactions 52. In the single-excitation subspace of the 1D system where the dephasing rate is much larger than energy relaxation, any initial state has a non-vanishing projection onto the equilibrium state $\rho_{E}=1/L\sum_{j}|1_{j}\rangle\langle 1_{j}|$ spreading in all sites of the lattice, which leads to a relaxation dynamics toward $\rho_{E}$ in a long-time limit $Rt\gg 1$ with a diffusion rate $R=2J^{2}/\Gamma$ 29. Mpemba-effect-like dynamics with dephasing noise is evaluated by the observable distance function $D(t)$ for a given state $\rho(t)$ from the equilibrium state $\rho_{E}$ defined as $D(t)=\operatorname{Tr}\left[\rho(t)\log\rho(t)\right]-\operatorname{Tr}\left[\rho_{E}(t)\log\rho_{E}(t)\right]$. In the case of single-particle excitation, it takes the following form in terms of the population $n_{j}(t)$ 30
$D(t)=\log L+\sum_{j}n_{j}(t)\log n_{j}(t).$ (3)
To experimentally observe the Mpemba-effect-like dynamics, a small system size is more favorable 30. As the system size $L$ increases, the uniformly distributed excitation of the equilibrium state has a smaller population at each site, and the relaxation dynamics become slower to reach equilibrium.
We now demonstrate the Mpemba-effect-like dynamics in our 1D chain of $L=7$ qubits, with the homogeneous coupling strengths $g_{j}=J\approx 2\pi\times 8.3$ MHz and $\Gamma/J\approx 3$ used to accelerate the diffusive dynamics toward equilibrium state $\rho_{E}$. As shown in Fig. 4(a), when the system is initially prepared as a highly localized pure state $\rho_{1}=|1_{j=0}\rangle\langle 1_{j=0}|$ with the excitation at site $j=0$, it shows spreading of qubit population $n_{j}(t)$ in all sites when the evolution time $t$ is large enough. In a similar scenario, we initialize the system as a mixed state $\rho_{4}=1/4\sum_{j=0}^{3}|1_{j}\rangle\langle 1_{j}|$ by randomly applying the $\pi$ pulse on the sites $j=0,1,2,3$, which is closer to the stationary state $\rho_{E}$ and gives a temporal evolution in Fig. 4(b), showing a slower diffusion toward the missing occupation on the other sites. Intuitively, the initial state with a closer distance to the equilibrium state is expected to diffuse faster and fill the missing occupation of other sites. To perform a systematic study, we prepare two more initial states $\rho_{2}=1/2\sum_{j=2}^{3}|1_{j}\rangle\langle 1_{j}|$, $\rho_{3}=1/3\sum_{j=1}^{3}|1_{j}\rangle\langle 1_{j}|$ between $\rho_{1}$ and $\rho_{4}$ in terms of distance to $\rho_{E}$, and observe their dynamics towards equilibrium similar to Figs. 4(a) and 4(b), see Supplemental Material for details and numerical simulations. The deviation between the measured population and corresponding numerical results is mainly limited by two-level system defects and readout errors. By evaluating the distance function $D(t)$ from the measured population $n_{j}(t)$, we present the results of $D(t)$ for different initial states in Fig. 4(c), which experimentally demonstrate that the farthest-from equilibrium state $\rho_{1}$ relaxes faster than the other three mixed states being closer to the equilibrium state $\rho_{E}$, clearly illustrating the Mpemba-effect-like dynamics.
In this work, we experimentally demonstrate the diffusive dynamics assisted by controlled dephasing noise in a 1D array of superconducting qubits. By properly arranging the coupling strengths between neighboring qubits, we implement a quasiperiodic lattice and demonstrate that dephasing can give distinct dynamical behavior in this system 29. Moreover, we explore the Mpemba-effect-like dynamics in this 1D model by preparing different excitation distributions in the qubit array and observing that more localized initial states relax to a uniformly distributed mixed state more rapidly under dephasing noise 30; 53. The experimental demonstration of these two counterexamples not only deepens our understanding of diffusive dynamics at the microscopic level, but also highlights the expanded potential of controlled dissipative processes in quantum simulation. Utilizing the dissipative effect synthesized in such fully controlled quantum systems, further efforts can be made for the studies, including exotic non-equilibrium phases 54; 55, reservoir computing approach 56, and quantum thermodynamics 57; 58.
See the supplementary material for detailed results on the device information, numerical simulations, and additional experiment results.
We thank Yucheng Wang, Xinchi Zhou, and Yuanzhen Chen for their helpful discussions. This work was supported by the Science, Technology and Innovation Commission of Shenzhen Municipality (KQTD20210811090049034, RCBS20231211090824040, RCBS20231211090815032), the National Natural Science Foundation of China (12174178, 12204228, 12374474 and 123b2071), the Innovation Program for Quantum Science and Technology (2021ZD0301703), the Shenzhen-Hong Kong Cooperation Zone for Technology and Innovation (HZQB-KCZYB-2020050), and Guangdong Basic and Applied Basic Research Foundation (2024A1515011714, 2022A1515110615).
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Author contributions
Y.L. and C.X. contributed equally to this work. C.X. and Y.L. performed the experiments and theoretical analysis under the supervision of W.R. and Z.T. Youpeng Z., W.R., and Z.T. supervised and coordinated the project. All authors contributed to the discussion and preparation of the manuscript.
Data availability
The data that support the findings of this study are available from the corresponding authors upon reasonable request.